# Generating multipartite nonlocality to benchmark quantum computers

### v2

@@ V Simulation for an IBM quantum computer
Figure 10: Average p-value of the results for the LC and GHZ state in Fig. 9. The average is taken over the p-values of the $10$ repetitions. The error bars denote the standard deviation. We plot only the top error bars as we are interested in the uncertainty to larger $p$-values.
Finally, we simulate the Bell inequalities of the LC and the GHZ state for the IBM Eagle quantum processor. For this purpose, we use the Qiskit AerSimulator [52] with the noise data of the quantum computer ibm_brisbane available at [45]. In Fig. 6, we show the ideal circuits to prepare the LC and the GHZ state for $n=6$ qubits. The advantage of the LC state is that it can be prepared by a circuit of constant depth of three, whereas the step complexity for the GHZ state increases with the number of qubits $n$. We note, however, that IBM’s Eagle processor does not implement the Hadamard and CZ gate natively. Rather, the gates have to be composed in terms of the available gate set. In practice the circuits thus involve more gates and exhibit a larger depth.
After the preparation, $L$ random terms of the corresponding Bell inequality are measured. The measurement of each random term is not repeated, i.e., $K=1$. Fig. 9 shows the average expectation values of the Bell inequalities for the LC and GHZ state of up to $n=24$ qubits. We have chosen $L=800$ random terms and the average is taken over $10$ repetitions. Moreover, the number of qubits is a multiple of three as only in this case a good Bell inequality for the LC state is known. Fig. 9 shows that for both states the simulation predicts a Bell violation that increases exponentially with $n$. The LC state however shows a slightly higher relative violation compared to the GHZ state. This can be seen in Fig. 11a. The plot in Fig. 11a displays the observed expectation value as a fraction of the quantum bound, i.e., ^$\langle{\hat{\mathcal{B}}}\rangle/Q$. In agreement with the scaling that is predicted by the depolarisation noise model, we fit the logarithmic data to a linear function for the LC state and a quadratic function for the GHZ state. The fits yield
@@
This affirms that the relative violation of the GHZ state decreases faster with $n$ compared to the LC state. We attribute this to the larger circuit depth that is required for the GHZ state. The preparation of the GHZ state is thus more affected by noise. The smaller relative violation is also the reason for the larger $p$-values for the GHZ state, which we present in Fig. 10. Except for the cases of $n=3,15$ the $p$-values of the LC state are smaller, which implies a higher significance of the observed violation. Overall, the $p$-values show a large variation. We note that the $p$-values depend on the observed violation. Apparently, for some numbers of qubits $n$ the evaluation of the Bell inequalities is more affected by noise. We attribute this to the mapping to the actual architecture. For different sizes of the state, the optimal choice of qubits might be differently affected by noise. This also explains the variation of the error bars in Fig.  9. In case the evaluation is more affected by noise, not only is the observed violation smaller but it also appears reasonable that the value fluctuates more. In addition, we note that the error bars in Fig. 11a do not match the fitted exponential. This is because the plotted error bars show the standard deviation of the repeated simulations and do not cover all uncertainties. For example, the error bars do not include the uncertainties due to the different mappings of the quantum circuits to the actual architecture.
Finally, we use Eq. (23) to extrapolate the violation to larger $n$. We note that the circuits have to be adapted to the architecture of the quantum computer. For this task, we use the automatic transpilation provided in Qiskit, which involves an optimization such that the qubits are chosen that are least affected of noise. The extrapolation overestimates the violation for large $n$, since in this case such a choice is not possible anymore. Hoeffding’s inequality in Eq. (17) yields for $K=1$ and a target value for the confidence $\gamma$:
$L(n,\gamma)\geq\left\lceil-\frac{2}{(t^{2}/M^{2})(n)}\log(1-\gamma)\right\rceil.$ (24)
The number of necessary sampled terms $L$ is shown in Fig. 11b for the LC state and in Fig. 11c for the GHZ state.
@@
Figure 11: (a) Observed violation as a fraction of the quantum bound, i.e., ^$\langle{\hat{\mathcal{B}}}\rangle/Q$. The logarithmic data is fitted by a linear function for the LC state and a quadratic function for the GHZ state. From the fitted function, we estimate the number of random terms $L$ that have to be sampled to reach a confidence of $\gamma=1\sigma,2\sigma,5\sigma$. (b) shows the results for the LC state, whereas (c) illustrates the predictions for the GHZ state.
@@
As an example, we discuss $L$ for the case of $n=108$ qubit LC state, i.e., the largest simple path for the architecture. For better comparison, we also calculate $L$ for the $n=108$ qubit GHZ state. To reach a target confidence of $\gamma=5\sigma$ the values are:
@@
Nonlocality of the LC state can be verified by a large number of measurement settings. For the GHZ state, in turn, the extrapolation predicts also a violation. To verify this violation, however, requires an infeasible number of measurements. This shows again that the GHZ state is more affected by noise than the LC state and that the LC state is thus more promising to detect large multipartite nonlocality. The above numbers are still smaller than $M_{\text{LC}}\approx 10^{21}$ and $M_{\text{GHZ}}\approx 10^{32}$, but much larger than the values predicted in Sec. IV. We attribute this to the fact that the generic depolarisation noise model does not cover all the noise in a specific quantum computer. Recall, however, that our main goal is to detect large-scale quantum nonlocality and our method can still be used to assess the confidence of an observed Bell violation. Of course, the verification with a certain confidence may require more measurements for specific computers. In the worst case (e.g., when $\alpha$ drops and the $p$-value is not sufficient), one just needs to make additional measurements.
@@ VI Conclusion
We have demonstrated a method to test $n$-partite Bell nonlocality on quantum computers. Quantum computer often have restricted two-qubit connectivity. We have thus pointed out that graph states are a natural choice of non-local states that can be readily prepared if the graph is a subgraph of the connectivity graph. Moreover, for certain graph states good Bell inequalities are known. These Bell inequalities allow for an exponential violation of the classical bound, but, in turn, also typically require an exponential number of measurements. On the one hand, the exponential violation makes them increasingly robust to noise. On the other hand, it is impossible to measure all terms in an experiment. We have solved this problem by proposing a method in the manner of randomized measurements, e.g., direct fidelity estimation [35, 36] or few-copy entanglement detection [37]. By sampling the terms of the Bell operator at random the number of measurements can be drastically reduced. The violation can, however, still be verified with high significance. We have gauged the significance of a result with Hoeffding’s inequality. It thus dependents on the violation observed. To assess the usefulness of the method on real devices, we have first used a simple depolarisation noise model to estimate realistic violations. Finally, we have simulated the method for the IBM Eagle quantum processor. As expected with increasing accuracy of the noise model the predicted violation shrinks. However, also the simulator of the IBM processor predicts the number of terms that have to be sampled to be much smaller than the total number of terms in the Bell inequalities.
@@
Furthermore, the method could also be refined. For example, as the Bell inequalities only include stabilizers of the graph state, all observables commute. It might thus be feasible to find a (possibly very complicated) POVM to simultaneously measure all terms.
Moreover, it could be interesting to analysis the Bell inequalities with other statistical methods. Instead of the $p$-value, one might look at the Kullback–Leibler divergence that has been used to assess the statistical strength of Bell inequalities for few parties [40].
The relation to other benchmarks, e.g., the quantum volume [53, 54] or the layer fidelity [55], is also yet to be explored. Especially the layer fidelity can be measured by benchmarking a linear string of qubits of the quantum computer.
Note added: After submission of our article similar ideas have been discussed in [56].
@@ Acknowledgements
The authors would like to thank Lina Vandré, H. Chau Nguyen, Mariami Gachechiladze, Konrad Szymański, Ties Ohst, Kiara Hansenne, and Carlos de Gois for useful discussions and comments. This work has been supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, project numbers 447948357 and 440958198), the Sino-German Center for Research Promotion (Project M-0294), and the German Ministry of Education and Research (Project QuKuK, BMBF Grant No. 16KIS1618K). JLB acknowledges support from the House of Young Talents of the University of Siegen. AC is supported by the EU-funded project FoQaCiA Foundations of Quantum Computational Advantage and the MCINN/AEI (Project No. PID2020-113738GB-I00).
@@ A.1 Estimator in the infinite measurement limit
First, we assume that the expectation values can be inferred directly, i.e., that we can repeat the measurement of the operator infinite times. In this case, the estimator is given by Eq. (13). The expectation value has to be calculated with respect to the random variables $J_{l}$. With $\mathbb{E}[\langle{B_{J_{l}}}\rangle]=\sum_{j=1}^{M}p(J_{l}=j)\langle{B_{j}}\rangle=\frac{1}{M}\sum_{j=1}^{M}\langle{B_{j}}\rangle$ we obtain
@@ A.2 Estimator for finite repetitions
To calculate the expectation value of the estimator in Eq. (16), we note that both the measurement outcomes $b_{j}$ and the index $J$ of the terms are random variables. Hence, the expectation value of the estimator has to be taken over both the measurement outcomes as well as the random picking, i.e., over $J$. To evaluate the expectation value we can thus make use of the law of iterated expectation. That, is,
@@ B.2 Estimator for finite repetitions
As in the previous section, we can apply Hoeffding’s inequality to the estimator in Eq. (16). Also the estimator in Eq. (16) is a sum of independent random variables:
@@
Since the outcomes $b_{J_{l}}^{(k)}$ are obtained from different experimental runs, they are independent and thus are the random variables $Y_{kl}$. In addition, the outcomes can only take the values $b_{J_{l}}^{(k)}\in\{-1,1\}$. Therefore, we have that the random variables $Y_{kl}$ are bounded as: $-\frac{M}{KL}=a_{kl}\leq Y_{kl}\leq b_{kl}=\frac{M}{KL}$. Finally, we get from Hoeffding’s inequality
@@ Appendix C Preparation scheme for the LC state
We discuss a scheme to prepare a LC state with all qubits of a $n$-qubit quantum computer. This can be done by preparing all qubits in the $\ket{+}$ state and then applying CZ gates between some of them. A problem can be that the connectivity of the quantum computer does not allow to perform a specific gate between qubits $i$ and $j$ directly. The following lemma shows that this is not a fundamental problem.
@@
We give an explicit construction that is visualized in Fig. 12. The initial state is shown in Fig. 12a. We would like to perform a CZ gate between qubits $1$ and $m$, i.e., $\text{CZ}_{1m}$. The interaction topology, however, does not allow a direct coupling. Rather the qubits $1$ and $m$ are connected by the path of qubits $1,2,\ldots m$. Here, all the qubits $1,2,\ldots m$ should be in the $\ket{+}$ state, especially, it is important that no CZ gate has been applied to the qubits $2,\ldots m$ yet. In this case, we can apply the following scheme:
  1. 1.
    Connect qubit $2$ by performing the $\text{CZ}_{12}$ gate to generate the first $\text{PAIR}(1,2)$ (Fig. 12b).
  2. 2.
    While $k<m$ transform $\text{PAIR}(1,k)\rightarrow\text{PAIR}(1,k+1)$ by
@@
      Couple qubit $k$ and $k+1$ by $\text{CZ}_{k,k+1}$ (Fig. 12c).
@@
      Couple qubit $1$ and $k+1$ by a local complementation on qubit $k$, i.e., $\text{LC}_{k}$ (Fig. 12d).
@@
      Cancel the CZ gate between qubits $1$ and $k$ by performing $\text{LC}_{k+1}$ (Fig. 12e).
@@
      Cancel the CZ gate between qubits $k$ and $k+1$ by the controlled-Z $\text{CZ}_{k,k+1}$ (Fig. 12f).
This allows to decompose the $\text{CZ}_{1m}$ gate into a sequence with circuit depth $13(m-2)+$ 1. We note that step $2$ is only necessary for $m>1$ and requires three steps as the local complementations in (b) and (c) can be combined. ∎
Figure 13: Illustration of the scheme in Obs. 4 for an exemplary two-qubit connectivity of eight qubits. The dotted lines denote the CZ gates that can be performed, whereas the red lines indicate the graph state.
@@
On a quantum computer of $n$-qubits it is possible to prepare a $n$-qubit LC state with $\mathcal{O}(n)$ circuit depth.
@@
  1. 1.
    We start at a leaf and successively couple the adjacent qubits in direction of the root by CZ operations.
  2. 2.
    At a branch-off, check whether the other branch has already been covered. If all other branches have already been covered, we continue step $1$ in direction of the root. Otherwise, Lemma 1 can be used to couple the last qubit to a leaf in the uncovered branch. From there we can continue again with step $1$.
The scheme is shown for an exemplary two-qubit connectivity in Fig. 13. To investigate the circuit depth, we note that step $1$ and $2$ are executed alternately. We thus count the number of steps for each run. $k_{i}$ denotes the number of steps for the $i$-th execution of step $1$, whereas $l_{i}$ stands for the steps required for the $i$-th execution of step $2$. In step $1$, adjacent qubits are coupled consecutively by CZ gates. We thus have $\sum_{i}k_{i}<n$. In each step $2$, a qubit at distance $m_{i}$ is coupled and from Lemma 1 we know that $l_{i}=3(m_{i}-2)+1$. As each branch is only passed once, we have $\sum_{i}(m_{i}-2)\leq n$. Moreover, there are less than $n$ branch-offs, i.e., $\sum_{i}1\leq n$. Therefore, we obtain $\sum_{i}l_{i}=\sum_{i}3(m_{i}-2)+1\leq 4n$. The final circuit depth of the scheme is thus upper bounded by $\sum_{i}(k_{i}+l_{i})\leq n+4n=5n$. ∎

### v3

@@ V Simulation for an IBM quantum computer
Finally, we simulate the Bell inequalities of the LC and the GHZ states for the IBM Eagle quantum processor. For this purpose, we use the Qiskit AerSimulator [52] with the noise data of the quantum computer ibm_brisbane available at [45]. In Fig. 6, we show the ideal circuits to prepare the LC and the GHZ states for $n=6$ qubits. The advantage of the LC state is that it can be prepared by a circuit of constant depth of three, whereas the step complexity for the GHZ state increases with the number of qubits, $n$. We note, however, that IBM’s Eagle processor does not implement the Hadamard and CZ gates natively. Rather, the gates have to be composed in terms of the available gate set. In practice, the circuits thus involve more gates and exhibit a larger depth.
After the preparation, $L$ random terms of the corresponding Bell inequality are measured. The measurement of each random term is not repeated, i.e., $K=1$. Figure 9 shows the average expectation values of the Bell inequalities for the LC and GHZ states of up to $n=24$ qubits. We have chosen $L=800$ random terms and the average is taken over $10$ repetitions. Moreover, the number of qubits is a multiple of three as only in this case is a good Bell inequality for the LC state known. Figure 9 shows that for both states, the simulation predicts a Bell violation that increases exponentially with $n$. The LC state, however, shows a slightly higher relative violation compared to the GHZ state. This can be seen in Fig. 11(a). The plot in Fig. 11(a) displays the observed expectation value as a fraction of the quantum bound, i.e., ^$\langle{\hat{\mathcal{B}}}\rangle/Q$. In agreement with the scaling that is predicted by the depolarization noise model, we fit the logarithmic data to a linear function for the LC state and a quadratic function for the GHZ state. The fits yield
@@
This affirms that the relative violation of the GHZ state decreases faster with $n$ compared to the LC state. We attribute this to the larger circuit depth that is required for the GHZ state. The preparation of the GHZ state is thus more affected by noise. The smaller relative violation is also the reason for the larger $p$ values for the GHZ state, which we present in Fig. 10. Except for the cases of $n=3,15$, the $p$ values of the LC state are smaller, which implies a higher significance of the observed violation. Overall, the $p$ values show a large variation. We note that the $p$ values depend on the observed violation. Apparently, for some numbers of qubits, $n$, the evaluation of the Bell inequalities is more affected by noise. We attribute this to the mapping of the actual architecture. For different sizes of the state, the optimal choice of qubits might be differently affected by noise. This also explains the variation of the error bars in Fig.  9. In case the evaluation is more affected by noise, not only is the observed violation smaller but it also appears reasonable that the value fluctuates more. In addition, we note that the error bars in Fig. 11(a) do not match the fitted exponential. This is because the plotted error bars show the standard deviation of the repeated simulations and do not cover all uncertainties. For example, the error bars do not include the uncertainties due to the different mappings of the quantum circuits to the actual architecture.
Figure 10: Average $p$ value of the results for the LC and GHZ states in Fig. 9. The average is taken over the $p$ values of the $10$ repetitions. The error bars denote the standard deviation. We plot only the top error bars as we are interested in the uncertainty to larger $p$ values.
@@
Figure 11: (a) Observed violation as a fraction of the quantum bound, i.e., ^$\langle{\hat{\mathcal{B}}}\rangle/Q$. The logarithmic data are fitted by a linear function for the LC state and a quadratic function for the GHZ state. From the fitted function, we estimate the number of random terms, $L$, that have to be sampled to reach a confidence of $\gamma=1\sigma,2\sigma,5\sigma$. (b) The results for the LC state; (c) the predictions for the GHZ state.
Finally, we use Eq. (23) to extrapolate the violation to larger $n$. We note that the circuits have to be adapted to the architecture of the quantum computer. For this task, we use the automatic transpilation provided in Qiskit, which involves an optimization such that the qubits are chosen that are least affected by noise. The extrapolation overestimates the violation for large $n$, since in this case such a choice is no longer possible. Hoeffding’s inequality in Eq. (17) yields the following for $K=1$ and a target value for the confidence $\gamma$:
$L(n,\gamma)\geq\left\lceil-\frac{2}{(t^{2}/M^{2})(n)}\ln(1-\gamma)\right\rceil.$ (24)
The number of necessary sampled terms, $L$, is shown in Fig. 11(b) for the LC state and in Fig. 11(c) for the GHZ state.
@@
As an example, we discuss $L$ for the case of the $n=108$-qubit LC state, i.e., the largest simple path for the architecture. For better comparison, we also calculate $L$ for the $n=108$-qubit GHZ state. To reach a target confidence of $\gamma=5\sigma$, the values are
@@
The nonlocality of the LC state can be verified by a large number of measurement settings. For the GHZ state, in turn, the extrapolation also predicts a violation. To verify this violation, however, requires an infeasible number of measurements. This shows again that the GHZ state is more affected by noise than the LC state and that the LC state is thus more promising to detect large multipartite nonlocality. The above numbers are still smaller than $M_{\text{LC}}\approx 10^{21}$ and $M_{\text{GHZ}}\approx 10^{32}$, but much larger than the values predicted in Sec. IV. We attribute this to the fact that the generic depolarization noise model does not cover all the noise in a specific quantum computer. Recall, however, that our main goal is to detect large-scale quantum nonlocality and our method can still be used to assess the confidence of an observed Bell violation. Of course, the verification with a certain confidence may require more measurements for specific computers. In the worst case (e.g., when $\alpha$ drops and the $p$ value is not sufficient), one just needs to make additional measurements.
@@ VI Conclusion
We have demonstrated a method to test $n$-partite Bell nonlocality on quantum computers. Quantum computers often have restricted two-qubit connectivity. We have thus pointed out that graph states are a natural choice of nonlocal states that can be readily prepared if the graph is a subgraph of the connectivity graph. Moreover, for certain graph states, good Bell inequalities are known. These Bell inequalities allow for an exponential violation of the classical bound, but, in turn, also typically require an exponential number of measurements. On the one hand, the exponential violation makes them increasingly robust to noise. On the other hand, it is impossible to measure all terms in an experiment. We have solved this problem by proposing a method in the manner of randomized measurements, e.g., direct fidelity estimation [35, 36] or few-copy entanglement detection [37]. By sampling the terms of the Bell operator at random, the number of measurements can be drastically reduced. The violation can, however, still be verified with high significance. We have gauged the significance of a result with Hoeffding’s inequality. It thus dependents on the violation that is observed. To assess the usefulness of the method on real devices, we have first used a simple depolarization noise model to estimate realistic violations. Finally, we have simulated the method for the IBM Eagle quantum processor. As expected with increasing accuracy of the noise model, the predicted violation shrinks. However, also, the simulator of the IBM processor predicts the number of terms that have to be sampled to be much smaller than the total number of terms in the Bell inequalities.
@@
Furthermore, the method could also be refined. For example, as the Bell inequalities only include stabilizers of the graph state, all observables commute. It might thus be feasible to find a (possibly very complicated) positive operator valued measure (POVM) to simultaneously measure all of the terms.
Moreover, it could be interesting to analyze the Bell inequalities with other statistical methods. Instead of the $p$ value, one might look at the Kullback-Leibler divergence that has been used to assess the statistical strength of Bell inequalities for few parties [40].
The relation to other benchmarks, e.g., the quantum volume [53, 54] or the layer fidelity [55], is also yet to be explored. In particular, the layer fidelity can be measured by benchmarking a linear string of qubits of the quantum computer.
Note added. Recently, similar ideas have been discussed in [56].
@@ Acknowledgements
The authors would like to thank Lina Vandré, H. Chau Nguyen, Mariami Gachechiladze, Konrad Szymański, Ties Ohst, Kiara Hansenne, and Carlos de Gois for useful discussions and comments. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team. This work has been supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, Projects No. 447948357 and No. 440958198), the Sino-German Center for Research Promotion (Project No. M-0294), and the German Ministry of Education and Research (Project QuKuK, BMBF Grant No. 16KIS1618K). J.L.B. acknowledges support from the House of Young Talents of the University of Siegen. A.C. is supported by the EU-funded project FoQaCiA Foundations of Quantum Computational Advantage and the MCINN/AEI (Project No. PID2020-113738GB-I00).
@@ A.1 Estimator in the infinite measurement limit
First, we assume that the expectation values can be inferred directly, i.e., that we can repeat the measurement of the operator infinite times. In this case, the estimator is given by Eq. (13). The expectation value has to be calculated with respect to the random variables $J_{l}$. With $\mathbb{E}[\langle{B_{J_{l}}}\rangle]=\sum_{j=1}^{M}p(J_{l}=j)\langle{B_{j}}\rangle=\frac{1}{M}\sum_{j=1}^{M}\langle{B_{j}}\rangle$, we obtain
@@ A.2 Estimator for finite repetitions
To calculate the expectation value of the estimator in Eq. (16), we note that both the measurement outcomes $b_{j}$ and the index $J$ of the terms are random variables. Hence, the expectation value of the estimator has to be taken over both the measurement outcomes as well as the random picking, i.e., over $J$. To evaluate the expectation value, we can thus make use of the law of iterated expectation. That is,
@@ B.2 Estimator for finite repetitions
As in the previous section, we can apply Hoeffding’s inequality to the estimator in Eq. (16). Also the estimator in Eq. (16) is a sum of independent random variables,
@@
Since the outcomes $b_{J_{l}}^{(k)}$ are obtained from different experimental runs, they are independent and thus are the random variables $Y_{kl}$. In addition, the outcomes can only take the values $b_{J_{l}}^{(k)}\in\{-1,1\}$. Therefore, we have that the random variables $Y_{kl}$ are bounded as $-\frac{M}{KL}=a_{kl}\leq Y_{kl}\leq b_{kl}=\frac{M}{KL}$. Finally, we get from Hoeffding’s inequality,
@@ Appendix C Preparation scheme for the LC state
We discuss a scheme to prepare a LC state with all qubits of an $n$-qubit quantum computer. This can be done by preparing all qubits in the $\ket{+}$ state and then applying CZ gates between some of them. A problem can be that the connectivity of the quantum computer does not allow one to perform a specific gate between qubits $i$ and $j$ directly. The following lemma shows that this is not a fundamental problem.
@@
We give an explicit construction that is visualized in Fig. 12. The initial state is shown in Fig. 12(a). We would like to perform a CZ gate between qubits $1$ and $m$, i.e., $\text{CZ}_{1m}$. The interaction topology, however, does not allow a direct coupling. Rather, the qubits $1$ and $m$ are connected by the path of qubits $1,2,\ldots m$. Here, all the qubits $1,2,\ldots m$ should be in the $\ket{+}$ state; in particular, it is important that no CZ gate has been applied yet to the qubits $2,\ldots m$. In this case, we can apply the following scheme:
  1. (1)
    Connect qubit $2$ by performing the $\text{CZ}_{12}$ gate to generate the first $\text{PAIR}(1,2)$ [Fig. 12(b))].
  2. (2)
    While $k<m$ transform $\text{PAIR}(1,k)\rightarrow\text{PAIR}(1,k+1)$ by the following:
@@
      Couple qubit $k$ and $k+1$ by $\text{CZ}_{k,k+1}$ [Fig. 12(c)].
@@
      Couple qubit $1$ and $k+1$ by a local complementation on qubit $k$, i.e., $\text{LC}_{k}$ [Fig. 12(d)].
@@
      Cancel the CZ gate between qubits $1$ and $k$ by performing $\text{LC}_{k+1}$ [Fig. 12(e)].
@@
      Cancel the CZ gate between qubits $k$ and $k+1$ by the controlled-Z $\text{CZ}_{k,k+1}$ [Fig. 12(f)].
This allows one to decompose the $\text{CZ}_{1m}$ gate into a sequence with circuit depth $13(m-2)+$ 1. We note that step (2) is only necessary for $m>1$ and requires three steps as the local complementations in (b) and (c) can be combined. ∎
@@
On a quantum computer of $n$ qubits, it is possible to prepare an $n$-qubit LC state with $\mathcal{O}(n)$ circuit depth.
@@
  1. (1)
    We start at a leaf and successively couple the adjacent qubits in the direction of the root by CZ operations.
  2. (2)
    At a branch-off, check whether the other branch has already been covered. If all other branches have already been covered, we continue step (1) in the direction of the root. Otherwise, Lemma 1 can be used to couple the last qubit to a leaf in the uncovered branch. From there, we can continue again with step (1).
The scheme is shown for an exemplary two-qubit connectivity in Fig. 13. To investigate the circuit depth, we note that step (1) and (2) are executed alternately. We thus count the number of steps for each run. $k_{i}$ denotes the number of steps for the $i$th execution of step (1), whereas $l_{i}$ stands for the steps required for the $i$th execution of step (2). In step (1), adjacent qubits are consecutively coupled by CZ gates. We thus have $\sum_{i}k_{i}<n$. In each step (2), a qubit at distance $m_{i}$ is coupled and, from Lemma 1 we know that $l_{i}=3(m_{i}-2)+1$. As each branch is only passed once, we have $\sum_{i}(m_{i}-2)\leq n$. Moreover, there are less than $n$ branch-offs, i.e., $\sum_{i}1\leq n$. Therefore, we obtain $\sum_{i}l_{i}=\sum_{i}3(m_{i}-2)+1\leq 4n$. The final circuit depth of the scheme is thus upper bounded by $\sum_{i}(k_{i}+l_{i})\leq n+4n=5n$. ∎
Figure 13: Illustration of the scheme in Observation 4 for an exemplary two-qubit connectivity of eight qubits. The dotted lines denote the CZ gates that can be performed, whereas the red lines indicate the graph state.